OpenStudy (anonymous):

Hi. Could someone please explain to me how to Graph Exponential Functions, preferably this one EXAMPLE: 7Log(3x)=15 solve for x.

OpenStudy (anonymous):

I am assuming a common log, i.e., base 10. There is more than one approach, but I would probably isolate the log by dividing by 7, and then using the definition of log to convert to an exponential form.

OpenStudy (anonymous):

\[\log(3x)=\frac{ 15 }{ 7 }\]

OpenStudy (anonymous):

thats how you would graph it?

OpenStudy (anonymous):

To graph you would normally have two variables; in this case there is no second variable such as a y. I was working towards solving for x.

OpenStudy (anonymous):

could you explain to me step by step how to graph and solve it..... if thats not to much to ask please?

OpenStudy (anonymous):

let's work on solving for x first. Again, I am assuming that we have common log (base 10).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait I had a typo

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use \[y=\log_{b}x<=> b ^{y}=x \]in this case we have\[10^{15/7}=3x\]

OpenStudy (anonymous):

then just divide both sides by 3\[x=\frac{ 1 }{ 3 }*10^{15/7}\]

OpenStudy (anonymous):

ok, i see

OpenStudy (anonymous):

This is an exact solution for x, and we can use it graph it.

OpenStudy (anonymous):

so i just use that? how does that work?

OpenStudy (anonymous):

it graphs just like any other relation that has the variable y, but the axes are rotated 90-degrees. usually an exponential function of this form staring with "y =" would roughly look likebut...